OpenStudy (anonymous):

what is the slope of the line that passes through the given points? (2,13) and (6,11)

OpenStudy (anonymous):

You missed some parethesis

OpenStudy (anonymous):

\[m=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

So in in this case (x1, y1) is (2,13) and (x2-y2) is (6,11) Plug in your numbers, 11-13/6-2 = -2/4 = -1/2